In Newton's ring experiment, sodium light is used (A1=589 nm, 12=589.6nm). Radius of curvature of plano-convex lens R=100 nm.

Calculate the order number of the ring at which they fades away. Find out the distance from the point of contact from where the rings fades away.
1
Expert's answer
2020-04-02T10:56:33-0400

"R=100 cm"


Sharpness of the rings pattern is the worst when the maxima and minima intermingle.

For the maxima


"r^2=\frac{R\lambda(2k-1)}{2}"


For the minima


"r^2=k\lambda R"


So,


"k\lambda_1R=\frac{R\lambda_2(2k-1)}{2}\to k\lambda_1=\frac{\lambda_2(2k-1)}{2}" , "\lambda_1=\lambda,\lambda_2=\lambda_1+\Delta\lambda=\lambda+\Delta\lambda"


"k\lambda=\frac {(\lambda+\Delta\lambda)(2k-1)}{2}\to k \Delta \lambda \approx\frac{\lambda}{2}"


"k=\frac{\lambda}{2\Delta\lambda}=\frac{\lambda_1}{2(\lambda_2-\lambda_1)}=\frac{589}{2(589.6-589)}\approx491"


"l=\sqrt{R^2-(R-\frac{k\lambda}{2})^2}=\sqrt{1^2-(1-\frac{491\cdot589\cdot10^{-9}}{2})^2}\approx0.017m=17mm"